### v1

# Emotional Dimension Control in Language Model-Based Text-to-Speech: Spanning a Broad Spectrum of Human Emotions

Abstract
Current emotional text-to-speech (TTS) systems face challenges in mimicking a broad spectrum of human emotions due to the inherent complexity of emotions and limitations in emotional speech datasets and models. This paper proposes a TTS framework that facilitates control over pleasure, arousal, and dominance, and can synthesize a diversity of emotional styles without requiring any emotional speech data during TTS training. We train an emotional attribute predictor using only categorical labels from speech data, aligning with psychological research and incorporating anchored dimensionality reduction on self-supervised learning (SSL) features. The TTS framework converts text inputs into phonetic tokens via an autoregressive language model and uses pseudo-emotional dimensions to guide the parallel prediction of fine-grained acoustic details. Experiments conducted on the LibriTTS dataset demonstrate that our framework can synthesize speech with enhanced naturalness and a variety of emotional styles by effectively controlling emotional dimensions, even without the inclusion of any emotional speech during TTS training.
Index Terms:
Emotional text-to-speech, emotional dimension, language model
I Introduction
Imagine a world similar to the one depicted in the movie ‘Her’, where a dialogue system not only talks but connects with users emotionally [1]. Emotional text-to-speech (TTS) systems are bringing us closer to this reality [2]. Designed to generate speech from text that conveys human-like emotions, emotional TTS has seen significant improvements thanks to recent advancements in deep learning [3]. As a crucial technology for human-computer interaction, emotional TTS is paving the way for dialogue systems that can be as engaging and empathetic as the personal companion portrayed in ‘Her’ [4].
However, current emotional TTS systems still face challenges stemming from the complexity of human emotions and limitations in both data and models [2]. Psychological studies suggest that humans can experience around 34,000 different emotions [5], yet current systems struggle to capture this full range. Previous emotional TTS research has either relied on supervising with categorical emotion labels [6, 7, 8, 9] or transferring styles from reference speech [10, 11]. Both approaches are constrained by the limited styles present in emotional speech datasets, often resulting in synthesized voices that exhibit averaged or stereotypical emotional patterns, thus providing limited control over emotion rendering [12, 13].
The scarcity of emotional speech data presents significant challenges for developing emotional TTS systems [2], particularly those that utilize language models [14, 15]. With large-scale training data, recent language model (LM)-based TTS frameworks [16, 17, 18] show strong in-context learning capabilities, effectively rendering appropriate prosody for complex text inputs and exhibiting robust zero-shot speaker cloning. Unlike traditional TTS task, which primarily focuses on capturing a wide range of phonetic variations across speakers [19], emotional TTS aims to render diverse emotional styles, thus requiring a large and diverse emotional speech data with detailed annotations. Emotional TTS systems like CosyVoice [14] and EmoCtrl-TTS [15], built on state-of-the-art language models, rely on thousands of hours of emotional speech data, underscoring the complexity and cost of such data collection. However, publicly available emotional speech datasets are typically small [20] and inadequate for training a robust LM-based TTS system from scratch. A common approach is to fine-tune a pre-trained TTS model with a small emotional dataset [21], but this can lead to overfitting and limit generalization across emotions and speakers [22]. Instead of scaling emotional TTS with larger datasets to achieve state-of-the-art benchmarks, this paper integrates insights from emotion theory to enable emotional dimension control and leverages the strong in-context learning capabilities of language models to build a TTS system that spans a broad spectrum of human emotions, especially when emotional data is limited or unavailable.
In this paper, we propose an effective approach for TTS to generate diverse emotional styles by controlling a emotional dimension vector. Our proposed method does not require any emotional speech data during TTS training and demonstrates the capability to span a broad spectrum of human emotions in the synthesized voices. Our approach is particularly inspired by Russell’s studies on emotion theory [23], which suggest that the 3 independent dimensions of pleasure, arousal, and dominance adequately capture the complexity of emotional expressions and are consistent across different contexts and subjects. We leverage the definition of three dimensions denoting emotions, as reported in [23], as foundational anchors to build a model aimed at predicting these values. The TTS framework then leverages pseudo-emotional dimensions to generate fine-grained acoustic details in parallel, based on the phonetic units predicted by the autoregressive model. During inference, our proposed framework facilitate emotional dimension control to synthesize diverse emotional styles without requiring any emotional references. Our contributions are summarize as below:
  • We propose an LM-based TTS framework with emotional dimension control over pleasure, arousal, and dominance, employing a 2-stage LM training approach that effectively renders emotions without requiring any emotional training data;
  • Our experiments show that our proposed model can synthesize a wide range of emotional styles, effectively spanning a broad spectrum of human emotions;
  • We also demonstrate the effectiveness of proposed method in enhancing the naturalness and prosody consistency with the prompt speech, outperforming the baseline.
We introduce our proposed method in Section 2. In Section 3, we report our experimental results. Section 4 concludes the paper.
Figure 1: An overview of the proposed text-to-speech (TTS) framework with emotional dimension control, consisting of: (a) Emotional Dimension (ED) Predictor Training, and (b) TTS Flow. The ED predictor is pre-trained on an emotional speech dataset to map emotional features to dimension representations via anchored dimensionality reduction. It then guides the non-autoregressive language model (LM) to predict acoustic details. ‘P’, ‘A’, and ‘D’ denote ‘Pleasure’, ‘Arousal’, and ‘Dominance’, ‘G2P’ represents ‘Grapheme-to-Phoneme’.
II Proposed Method
Following our previous research [24], we consider TTS as a conditional language modeling task, where the speech waveform is quantized into discrete tokens and is then modeled with a language model. However, training language models typically requires a large amount of data to achieve stable performance [25, 26], making it impractical as to emotional speech. Speech emotion, as a form of speech expressiveness, can be represented through continuous emotional dimensions that are learnable from expressive speech [27]. Our goal is to develop a TTS model that learns from expressive speech data during training and synthesizes speech that conveys intended emotions. The overall TTS system includes an autoregressive (AR) decoder and a non-autoregressive (NAR) decoder. The AR decoder predicts phonetic tokens from phoneme inputs and prompt audio, while the NAR decoder, using emotional dimension (ED) embeddings, predicts acoustic codecs with fine-grained details. Finally, the Encodec decoder synthesizes the speech waveform from these codecs.
II-A Predicting Emotional Dimensions from Categories
Human emotions can be represented categorically [28] or dimensionally [29]. Compared to categorical emotional approach [28] that lacks the capability to capture subtle emotional differences, the dimensional approach, such as Russell’s circumplex model, views emotions as fluid experiences across several dimensions, offering a more nuanced understanding [29]. The specific definitions [23] of 151 emotion types, as rated by 300 subjects, were reliably measured on a pleasure-arousal-dominance scale as shown in Table I. Despite those studies, most emotional TTS research still relies on a finite set of emotion labels [30, 22]. Few works have explored synthesizing mixed emotional styles by combining basic emotions [31, 32]. Current TTS systems [33, 34, 15] that control emotional dimensions rely on SER models trained with costly, labor-intensive emotional dimension annotations, and face similar biases and challenges as those encountered in SER [35, 36].
Inspired by a recent SER study [37] predicting arousal-valence only from categorical emotions, we design an ED predictor to map discrete emotions into a pleasure-dominance-arousal space (Figure 1(a)). We train the ED predictor with an emotional speech dataset, and each emotion category is assigned an anchor based on [23] to guide training, with initial vectors perturbed by a Gaussian noise $\theta$ = 0.01. Using pre-trained WavLM [38] for feature extraction, we predict emotional labels through a linear and classification layer. We derive a 128-d emotional feature vector from the linear layer and fine-tune the ED vectors from their initial settings. We create a $k$-Nearest Neighbor (kNN) graph and integrate categorical labels to differentiate emotion classes. UMAP [39] refines the vectors by minimizing cross-entropy loss between high-dimensional features and ED vectors. The ED predictor then estimates pleasure-arousal-dominance values for new speech samples, guiding acoustic predictions in the TTS model.
TABLE I: Some examples of the definition [23] of emotional states in terms of pleasure (‘P’), arousal (‘A’), and dominance (‘D’).
Emotion (P, A, D) Emotion (P, A, D)
Angry (-0.51, 0.59, 0.25) Elated (0.50, 0.42, 0.23)
Happy (0.81, 0.51, 0.46) Alert (0.49, 0.57, 0.45)
Sad (-0.63, -0.27, -0.33) Protected (0.60, -0.22, -0.40)
Surprise (0.40, 0.67, -0.13) Relaxed (0.68, -0.46, 0.20)
Anxious (0.01, 0.59, -0.15) Neutral (0.00, 0.00, 0.00)
II-B TTS Training with Emotional Dimension Guidance
Since speech contains both linguistic and para-linguistic information, we decouple TTS training as two stages: ‘Phoneme $\rightarrow$ Phonetic’ and ‘Phonetic $\rightarrow$ Acoustic’. The overall TTS diagram is illustrated in Figure 1(b). In the first stage, the AR model learns to map phoneme sequences to phonetic tokens. We leverage WavLM features with K-means clustering as phonetically-rich targets to train the AR model. In the second stage, the TTS focus on acoustic modeling with emotional dimension guidance, where a NAR LM is trained to recover fine-grained acoustic details from phonetic tokens predicted by AR model. Emotional dimension vectors, derived from the pre-trained ED predictor, are used to condition the NAR model, together with predicted phonetic tokens and acoustic tokens generated by Encodec. During TTS training, the AR model solely focuses on linguistic modeling, while the NAR model learns fine-grained acoustic details from the prompt audio, emphasizing the consistency of pleasure-arousal-dominance with the prompt.
II-C Emotion Rendering with Emotional Dimension Control
During inference, given input text and prompt audio along with its text transcriptions, the AR model predicts phonetic tokens representing phonetic variations over the phonemes. Subsequently, the NAR model predicts acoustic tokens based on the emotional dimension vector, the acoustic prompt, and the phoneme sequences. The emotional dimension vector can either be predicted from the prompt audio (‘ED prediction’) or manually assigned by users (‘ED control’). Finally, the encodec decoder synthesizes the speech waveform based on the predicted acoustic codecs.
III Experiments
III-A Experimental Setup
The emotional dimension predictor is trained on the train set from the English subset of the ESD dataset [40], which includes around 10 hours of emotional speech data across 5 basic emotions: Neutral, Angry, Happy, Sad, and Surprise. The emotional dimension predictor reported an 81% emotion classification accuracy on the ESD test set. We train the TTS framework on LibriTTS dataset [41], which consists of around 600 hours of speech data with transcriptions at a 24kHz sampling rate from 2,456 speakers. We merge the ‘train-clean’ and ‘train-other’ subsets as the training set and evaluate TTS performance on the ‘test-clean’ subset. Note that while LibriTTS contains expressive speech, it does not include any emotion labels. There is no emotional speech included in the TTS training.
To set up the experiments, for each speaker in the test set of LibriTTS, we randomly choose one speech sample from the same speaker to serve as the acoustic prompt, distinct from the speech to be synthesized, and then synthesize the speech based on the transcriptions. We compare 2 different inference settings:
  • Proposed with ED prediction, where ED values are predicted from the prompt audio;
  • Proposed with ED control, where ED values are assigned by users. In our experiments, ED values are directly derived from the literature [23] to simulate 5 basic emotions [28]111 5 Basic Emotions: Happy (0.81, 0.51, 0.46), Surprise (0.40, 0.67, -0.13), Angry (-0.51, 0.59, 0.25), Sad (-0.63, -0.27, -0.33), Neutral (0, 0, 0).; or redefined from the range in [23] to simulate 4 emotion pairs with subtle emotional differences2224 Emotion Pairs: Elation (1, 1, 0.5), Pleasure (1, 0, 0.5), Relaxed (0.5, -0.5, 0.2), Protected (0.5, -0.2, -0.5), Alert (0.5, 0.5, 0.5), Surprise (0.5, 0.5, -0.1), Anger (-1, 1, 1), Anxiety (-1, 1, -1), aiming to study relationships between 3 dimensions under controlled conditions;
We select training speech data that is up to 10 seconds long, and all synthesized speech samples range from 3 to 10 seconds. Speech demos are publicly available online [^3] Speech Demos: https://kunzhou9646.github.io/emo-icassp25/.
III-B Implementation Details
The TTS framework consists of both AR and NAR models, both of which consist of a transformer architecture consisting of 12 layers, 16 attention heads, an embedding dimension of 1024, a feed-forward layer dimension of 4096 and a dropout rate of 0.1. The 3-D emotional dimension vector can be either extracted from the prompt audio or assigned by users, which is then projected by a linear layer of dimension 128. We use a 24kHz Encodec [^4] https://huggingface.co/facebook/encodec_24khz as the acoustic encoder to compress the speech waveform into 8 layers of discrete acoustic tokens, while the Encodec decoder is utilized to recover 8 layers of discrete acoustic tokens to the speech waveform. We obtain the phonetic tokens from the 24th layer of WavLM [^5] https://huggingface.co/microsoft/wavlm-large models and cluster them using the K-means algorithm [42]. We manually upsample the phonetic tokens from 16kHz to 24kHz to align with acoustic tokens. The WavLM model for emotional dimension prediction is fine-tuned for 100 epochs with a batch size of 64 using the Adam optimizer at a 1e-4 learning rate. Dimensionality reduction settings are a 0.1 minimum distance and 20 nearest neighbors, with a 1e-2 learning rate for point optimization.
(a) 5 basic emotions controlled by emotional dimension vector1
(b) 8 emotions controlled by emotional dimension vector2
Figure 2: Comparison of pitch, energy, and spectral flux statistics for emotions generated and controlled by the emotional dimension vector across 1,731 speech samples from the LibriTTS ‘test-clean’ set.
Figure 3: Visualization of classification probabilities for 1,731 synthesized speech from the LibriTTS ‘test-clean’, using the emotion2vec model [43], a pre-trained 9-class speech emotion recognition (SER).
III-C Objective Evaluation
We visualize the statistics of pitch, energy, and spectral flux for all synthesized results. These 3 objective metrics are widely used in the SER field, as exemplified by tools like OpenSMILE [44]. Higher pitch and energy levels often indicate ‘Happy’ or ‘Surprise’, while greater spectral flux reflects rapid changes, associated with intense emotions like ‘Excitement’. We compare 5 basic emotions, as shown in Figure 2(a). It is observed that ‘Happy’ and ‘Surprise’ exhibit higher pitch, energy, and spectral flux, while ‘Sad’ shows the lowest values among the 5 emotions. As for 4 emotion pairs shown in Figure 2(b), ‘Elation’, ‘Pleasure’, ‘Alert’, and ‘Surprise’ exhibit higher pitch, energy, and spectral flux, whereas ‘Relaxed’ and ‘Protected’ show the lowest values. These results show that our model can synthesize distinguishable emotions solely by adjusting the emotional dimensions, aligning with the expected emotional characterizations.
We further employ a publicly available pre-trained speech emotion recognizer (SER), Emotion2vec [^6] https://huggingface.co/emotion2vec/emotion2vec_base, to evaluate all synthesized results. Emotion2vec [43], trained on 9 categories, is used to visualize the classification probabilities for 5 basic emotions, as shown in Figure 3. For 3 out of 5 intended emotions, the corresponding classification probabilities are the highest, demonstrating that our synthesized emotions can be perceptually recognized by the SER.
(a) Mean opinion score (MOS) to evaluate speech naturalness
(b) XAB Preference to evaluate speech naturalness and prosody similarity with the prompt speech
Figure 4: Subjective evaluation results assess the zero-shot performance on the TTS task. In the proposed framework, the emotional dimension (ED) is predicted based on the prompt speech (‘Proposed w/ ED Prediction’).
TABLE II: Comparative Mean Opinion Score (CMOS) for assessing emotional intelligibility of 4 synthesized emotions (‘Angry’, ‘Happy’, ‘Sad’, ‘Surprise’) compared to synthesized neutral speech.
Emotion Angry Happy Sad Surprise
CMOS 0.64 $\pm$ 0.33 0.71 $\pm$ 0.25 0.15 $\pm$ 0.32 0.92 $\pm$ 0.38
Figure 5: XAB test to assess the emotion intelligibility for 4 emotion pairs (‘Surprise’ and ‘Alert’, ‘Relaxed’ and ‘Protected’, ‘Elation’ and ‘Pleasure’, ‘Anger’ and ‘Anxiety’).
III-D Subjective Evaluation
We conducted listening experiments, with 10 native English speakers each listening to a total of 260 synthesized speech samples. We first evaluate our proposed framework on the zero-shot TTS task, where the ED is predicted based on the prompt speech (‘Proposed w/ ED Prediction’). We selected ‘Phonetic-Enhanced VALL-E’ [24] as the baseline. Compared to the popular VALL-E model, this baseline offers greater robustness and enhanced zero-shot capabilities [24]. A mean opinion score (MOS) test is conducted to evaluate the speech naturalness, where participants were asked to score each sample on a 5-point scale (5: Excellent, 4: Good, 3: Satisfactory, 2: Unsatisfactory, 1: Bad). As shown in Fig. 4(a), our proposed with ED prediction achieves a score of 3.67, outperforming the baseline score of 2.87.
We additionally conducted two AB preference tests, where participants were asked to choose the better speech sample in terms of speech naturalness and prosody similarity with the prompt speech, respectively. As shown in Fig. 4(b), the proposed framework with ED prediction consistently outperforms the baseline by achieving a higher preference percentage 78.9%. These results demonstrate that our proposed method significantly enhances speech naturalness and prosody consistency with the prompt by using ED to guide acoustic prediction during TTS training.
Next, we perform emotion perception experiments. Note that participants were not provided with explicit emotional references and they were asked to make judgments based on their innate emotional concepts. This approach aligns with methodologies used in emotion perception studies [45] and speech synthesis evaluation [31]. We first conducted a Comparative MOS (CMOS) test towards basic emotions1. In the proposed framework, the ED was manually assigned to simulate these emotions. In this test, participants were presented with pairs of synthesized neutral and emotional speech and asked to rate each sample on a 5-point scale based on how easily they could identify the intended emotion (5: Very Easy, 4: Easy, 3: Fair, 2: Difficult, 1: Very Difficult). The final results in Table II are measured as the comparative average score (CMOS) of emotional speech relative to the neutral counterpart. We observe that all emotions consistently achieve positive scores, indicating that listeners could perceive the intended emotions over the neutral one. Notably, ‘Surprise’ achieved the highest CMOS 0.92, demonstrating greater intelligibility, while ‘Sad’ had the lowest score 0.15, indicating it is closer to a neutral emotion.
We further conducted four XAB tests, grouping 4 emotion pairs 2 : 1) ‘Alert’ (0.5, 0.5, 0.5) vs. ‘Surprise’ (0.5, 0.5, -0.1); 3) ‘Relaxed’ (0.5, -0.5, 0.2) vs. ‘Protected’ (0.5, -0.2, -0.5); 3) ‘Anger’ (-1, 1, 1) vs. ‘Anxiety’ (-1, 1, -1); 4) ‘Elation’ (1, 1, 0.5) vs. ‘Pleasure’ (1, 0, 0.5). For each pair, participants were asked to correctly match the emotions, such as choosing between option A: X is Anger and Y is Anxiety and option B: X is Anxiety and Y is Anger. The results are summarized in Fig. 5. For all 4 emotion pairs, the correct match consistently exceeds 50%, with the ‘Elation vs. Pleasure’ pair achieving the highest accuracy at around 71.1%. The results in Fig.5 demonstrate a positive alignment with 2 key findings in [23]:
  • ‘Dominance’ is a key factor in distinguishing ‘Angry’ from ‘Anxious’, ‘Alert’ from ‘Surprised’, and ‘Relaxed’ from ‘Protected’ [^7] The first word in each pair involves dominance; the second word involves submissiveness.;
  • ‘Pleasure’ and ‘Arousal’ combine to contribute to ‘Elation’;
These observations indicate that our proposed TTS system can render emotions in accordance with psychological theories on the multidimensional nature of emotions, incorporating factors such as pleasure, arousal, and dominance. This enables our model to synthesize speech with a variety of emotional styles, enhancing both the naturalness and emotional intelligibility, and effectively bridging psychological theory with speech synthesis applications.
IV Conclusion
Inspired by the success of language models in a variety of generative tasks, much of current research prioritizes data collection, aiming to expand training data to push the boundaries of the state-of-the-art. In contrast, this paper explores how earlier studies on emotion theory can can advance research in emotional TTS. We introduce a TTS framework capable of controlling 3 emotional dimensions—pleasure, arousal, and dominance—during runtime. This framework not only synthesizes a wide range of emotional styles without the need for any emotional speech data for TTS training, but also enhances naturalness and prosody consistency with the speech prompt in zero-shot scenarios.