### v2

# Zhou Zhang Zhao Wang Pan Ng Zhang Ni Ma Nguyen Yip Ma Alibaba GroupSingapore RochesterUnited States of America

Abstract
Current emotional text-to-speech (TTS) systems face challenges in conveying the full spectrum of human emotions, largely due to the inherent complexity of human emotions and the limited range of emotional labels in existing speech datasets. To address these limitations, this paper introduces a TTS framework that provides flexible user control over three emotional dimensions—pleasure, arousal, and dominance—enabling the synthesis of a diverse array of emotional styles. The framework leverages an emotional dimension predictor, trained solely on categorical labels from speech data and grounded in earlier psychological research, which is seamlessly integrated into a language model-based TTS system. Experimental results demonstrate that the proposed framework effectively learns emotional styles from expressive speech, eliminating the need for explicit emotion labels during TTS training, while enhancing the naturalness and diversity of synthesized emotional speech.
keywords:
Emotional text-to-speech, emotional dimension, language model
1 Introduction
Emotional text-to-speech (TTS) aims to generate speech from text that conveys human-like emotions [1]. Recent advances in deep learning have significantly enhanced its capabilities. As a pivotal technology for human-computer interaction, emotional TTS facilitates the development of dialogue systems that are both engaging and empathetic [2].
However, current emotional TTS systems face challenges stemming from the complexity of human emotions and the limited range of emotional labels available in existing databases [2]. Psychological research suggests that humans can experience around 34,000 distinct emotions [3], yet current systems struggle to capture this full spectrum. Previous emotional TTS studies have primarily relied on two approaches: supervision using categorical emotion labels [4, 5] and style transfer from reference speech [6, 7]. Both approaches are constrained by the limited diversity of emotional styles present in available datasets, often resulting in synthesized speech with averaged or stereotypical emotional expressions, which restricts fine-grained control over emotion rendering [8, 9].
Recent language model (LM)-based TTS frameworks [10, 11] show strong in-context learning capabilities. With sufficient training data, these models effectively render appropriate prosody for complex text inputs and exhibit robust zero-shot speaker cloning capabilities. Unlike traditional TTS task, which focus primarily on capturing a wide range of phonetic variations across speakers [12], emotional TTS emphasizes rendering diverse emotional styles, thus requiring large and varied emotional speech datasets with detailed emotion annotations. Emotional TTS systems like CosyVoice [13] and EmoCtrl-TTS [14], built on state-of-the-art language models, require thousands of hours of emotional speech data, highlighting the complexity and cost of data collection. However, publicly available emotional speech datasets are typically small [15], making them inadequate for training robust LM-based TTS system from scratch. A common strategy is to fine-tune a pre-trained TTS model using small emotional datasets [16], but this approach risks overfitting and limits generalization across emotions and speakers [17]. To address these limitations, we propose a novel framework that synthesizes more nuanced and subtle emotions not present in the dataset by controlling continuous emotional dimensions. This method offers greater flexibility for emotion rendering and leverages the in-context learning capabilities of language models, enabling the generation of a wide spectrum of human emotions, even with limited or no emotion-labeled data.
In this paper, we propose an effective approach for LM-based TTS systems to generate diverse emotional styles by controlling an emotional dimension vector. Our method draws inspiration from Russell’s emotion theory [18], which identifies three independent dimensions—pleasure, arousal, and dominance—as sufficient to capture the complexity of emotional expressions across various contexts and individuals. Using these dimensions as foundational anchors, we develop a model that predicts their values. The pre-trained dimension predictor is then integrated into a LM-based TTS system to assist in next-token prediction, alongside the speaker embedding. During inference, our framework enables direct control over emotional dimensions, allowing the synthesis of diverse emotional styles without requiring emotional reference inputs. Our contributions are summarize as below:
  • We introduce a novel approach for controlling emotional dimensions in language model-based TTS systems, enabling fine-grained emotional speech synthesis;
  • Our approach learns emotional styles directly from expressive speech, eliminating the need for explicit emotion labels during TTS training;
  • During inference, the model allows users to control and explore a broad spectrum of human emotions by adjusting the emotional dimension vector. Grounded in established psychological theory, this approach ensures that the synthesized emotional styles align with the full range of human emotions;
  • We further demonstrate that our method enhances naturalness in zero-shot speaker cloning tasks, consistently outperforming the baseline system;
We introduce our proposed method in Section 2. In Section 3, we report our experimental results. Section 4 concludes the paper.
Figure 1: An overview of the proposed text-to-speech (TTS) framework with emotional dimension control, consisting of: (a) Emotional Dimension (ED) Predictor Training, and (b) Text-to-Speech Flow. The ED predictor is pre-trained on an emotional speech dataset to map emotional features to dimension representations via anchored dimensionality reduction. It then guides the autoregressive language model (LM) to predict acoustic details. ‘P’, ‘A’, and ‘D’ denote ‘Pleasure’, ‘Arousal’, and ‘Dominance’.
2 Proposed Method
2.1 Problem Formulation
We model TTS as a conditional language modeling task, where the speech waveform is quantized into discrete tokens and is then modeled with a language model [10, 13, 12]. However, training language models typically requires a large amount of data to achieve stable performance [19, 20], making it impractical as to emotional speech. Speech emotion, as a form of speech expressiveness, can be represented through continuous emotional dimensions that are learnable from expressive speech [21]. Our goal is to develop a TTS model that learns from expressive speech data during training and synthesizes speech that conveys intended emotions. Similar with that of CosyVoice [13], the overall TTS system includes an autoregressive LM decoder and a flow-matching module. The autoregressive LM decoder predicts acoustic tokens from text inputs and prompt speech using emotional dimension (ED) embeddings. Subsequently, a flow-matching module predicts the Mel-spectrum from the discrete acoustic tokens. Finally, the HiFi-GAN vocoder synthesizes the speech waveform from the Mel-spectrum.
2.2 Predicting Emotional Dimensions from Categories
Human emotions can be represented categorically [22] or dimensionally [23]. Compared to categorical emotional approach [22] that lacks the capability to capture subtle emotional differences, the dimensional approach, such as Russell’s circumplex model, views emotions as fluid experiences across several dimensions, offering a more nuanced understanding [23]. The specific definitions [18] of 151 emotion types, as rated by 300 subjects, were reliably measured on a pleasure-arousal-dominance scale as shown in Table 1. Despite those studies, most emotional TTS research still relies on a finite set of emotion labels [24, 17]. Few works have explored synthesizing mixed emotional styles by combining basic emotions [25, 26]. Current TTS systems [27, 28, 14] that control emotional dimensions rely on SER models trained with costly, labor-intensive emotional dimension annotations, and face similar biases and challenges as those encountered in SER [29, 30].
Inspired by a recent SER study [31] predicting arousal-valence only from categorical emotions, we design an ED predictor to map discrete emotions into a pleasure-dominance-arousal space (Figure 1(a)). We train the ED predictor with an emotional speech dataset, and each emotion category is assigned an anchor based on [18] to guide training, with initial vectors perturbed by a Gaussian noise $\theta$ = 0.01. Using pre-trained WavLM [32] for feature extraction, we predict emotional labels through a linear and classification layer. We derive a 128-d emotional feature vector from the linear layer and fine-tune the ED vectors from their initial settings. We create a $k$-Nearest Neighbor (kNN) graph and integrate categorical labels to differentiate emotion classes. UMAP [33] refines the vectors by minimizing cross-entropy loss between high-dimensional features and ED vectors. The ED predictor then estimates pleasure-arousal-dominance values for new speech samples, guiding acoustic predictions in the TTS model.
Table 1: Examples of the definition [18] of emotional states in terms of pleasure (‘P’), arousal (‘A’), and dominance (‘D’).
Emotion P A D Emotion P A D
Angry $-0.51$ $0.59$ $0.25$ Excited $0.62$ $0.75$ $0.38$
Happy $0.81$ $0.51$ $0.46$ Alert $0.49$ $0.57$ $0.45$
Sad $-0.63$ $-0.27$ $-0.33$ Protected $0.60$ $-0.22$ $-0.40$
Surprise $0.40$ $0.67$ $-0.13$ Relaxed $0.68$ $-0.46$ $0.20$
Anxious $0.01$ $0.59$ $-0.15$ Neutral $0.00$ $0.00$ $0.00$
2.3 TTS Training with Emotional Dimension Guidance
Following CosyVoice [13], we decompose the TTS process into two stages: ‘text-to-token’ and ‘token-to-speech’. The overall TTS diagram is illustrated in Figure 1(b). Given prompt speech, emotional dimension vectors are extracted using the pre-trained ED predictor (Figure 1(a)), while the speaker embedding is obtained from a pre-trained voiceprint model [^1] https://github.com/alibaba-damo-academy/3D-Speaker/tree/main/egs/3dspeaker/sv-cam++. The input text is first processed through a Grapheme-to-Phoneme tokenizer and then mapped to a shared semantic space with speech tokens using a text encoder. These extracted features collectively serve as conditioning inputs for the speech tokenizer, which produces speech token sequences.
We then frame the TTS task as an auto-regressive speech token generation task using a language model (LM). During LM training, we employ a teacher-forcing scheme to optimize LM predictions, where the left-shifted sequence serves as the input, and the original sequence acts as the expected output. We compute the cross-entropy loss between the predicted and ground-truth speech tokens, ensuring effective learning of token dependencies and accurate reconstruction of speech sequences.
2.4  Emotion Rendering with Emotional Dimension Control
During inference, the emotional dimension vector can be obtained in two ways: either be predicted from the prompt speech (‘ED prediction’) or manually assigned by users (‘ED control’), allowing for flexible emotion customization. Given the input text and prompt speech, the language model (LM) autoregressively predicts a sequence of speech tokens, which encode both linguistic and prosodic information. These tokens are then processed by a pre-trained optimal-transport flow-matching model, which learns the underlying distribution of Mel spectrograms. The flow-matching model generates high-fidelity Mel spectrogram samples that are conditioned on the predicted speech tokens, ensuring smooth and natural prosody. Finally, the synthesized Mel spectrograms are passed through a HiFi-GAN vocoder, which converts them into a high-quality speech waveform. Our model is expected to generate expressive and natural speech while preserving the desired emotional characteristics, whether inferred from the prompt speech or explicitly controlled by the user.
Figure 2: Statistical analysis of pitch, energy, and spectral flux for nine synthesized emotions, obtained by modulating the emotional dimension vector within the proposed framework using ED control (’ED Control’).
3 Experiments
Table 2: Mean Opinion Score (‘MOS’) of naturalness for the zero-shot TTS task. In the proposed system, emotional dimension values are predicted from the prompt speech (‘ED Prediction’).
Ground Truth CosyVoice [13] Proposed
4.80 $\pm$ 0.08 4.36 $\pm$ 0.13 4.54 $\pm$ 0.18
Figure 3: Mean Opinion Score for assessing emotional intelligibility (‘E-MOS’). In the proposed system, ED values were predicted based on the emotional reference (‘ED Prediction’).
3.1 Experimental Setup
The emotional dimension predictor is trained on the train set from the English subset of the ESD dataset [34], which includes around 10 hours of emotional speech data across 5 basic emotions: Neutral, Angry, Happy, Sad, and Surprise. The emotional dimension predictor reported an 81% emotion classification accuracy on the ESD test set. We train the TTS framework on LibriTTS dataset [35], which consists of around 600 hours of speech data with transcriptions at a 24kHz sampling rate from 2,456 speakers. We merge the ‘train-clean’ and ‘train-other’ subsets as the training set and evaluate TTS performance on the ‘test-clean’ subset. Note that while LibriTTS only contains expressive speech, it does not include any explicit emotion labels.
To set up the experiments, for each speaker in the test set of LibriTTS, we randomly choose one speech sample from the same speaker to serve as the acoustic prompt, distinct from the speech to be synthesized, and then synthesize the speech based on the transcriptions. We compare 2 different inference settings:
  • Proposed with ED prediction, where ED values are predicted from the prompt audio;
  • Proposed with ED control, where ED values are directly derived from the literature [18] to simulate different emotions, as shown in Table 1;
All synthesized speech samples range from 3 to 10 seconds. Speech demos are publicly available online [^2] Speech Demos: https://demos46.github.io/is25/.
3.2 Implementation Details
The backbone TTS framework was implemented based on the official release of CosyVoice [13], which consists of a speech tokenizer, a text encoder, a text-to-token LM and a flow-matching module. The speech tokenizer employs a ESPNet Conformer ASR model along with a vector quantizer of a codebook size of 4096 after the first 6 encoder layers. The text encoder consists of 6 transform layers with an 8-head and 512-d attention layer. The LM contains 12 transform layers with an 8-head and 512-d attention layer. An optimal-transport conditional flow matching model (OT-CFM) [36] is employed to predict the Mel spectrograms. A HiFi-GAN [^3] https://github.com/jik876/hifi-gan vocoder recovers the speech signal from the predicted Mel spectrogram.
The emotional dimension predictor consists of a WavLM features extractor followed by a linear layer, and an anchored dimensionality reduction module. The WavLM model is fine-tuned for 100 epochs with a batch size of 64 using the Adam optimizer at a 1e-4 learning rate. Dimensionality reduction settings are a 0.1 minimum distance and 20 nearest neighbors, with a 1e-2 learning rate for point optimization.
3.3 Objective Evaluation
We compute the statistics of pitch, energy and spectral flux for all synthesized speech, as shown in Figure 2. These acoustic features play a crucial role in characterizing emotional expressiveness. Higher pitch and energy levels are commonly associated with emotions such as ‘Excitement’, where increased vocal effort and higher frequency components convey heightened arousal. Similarly, greater spectral flux, which reflects rapid spectral variations, is indicative of dynamic speech patterns, often linked to intense emotions like ‘Surprise’, which exhibit abrupt prosodic changes.
From Figure 2, we observe that emotions such as ‘Excited’, ‘Surprise’, ‘Happy’ and ‘Alert’ exhibit relatively higher values across all three acoustic features, suggesting that these emotions are characterized by greater vocal dynamism and expressiveness. In contrast, emotions like ‘Sad’, ‘Angry’, ‘Neutral’ and ‘Protected’ tend to have lower values, indicating more subdued, steady, or restrained speech patterns. For instance, ‘Sad’ is typically marked by lower pitch and energy, corresponding to a softer, more monotonous voice, while ‘Neutral’ maintains moderate values across all features, aligning with a balanced, unexpressive tone. These findings provide quantitative evidence that our model successfully synthesizes distinguishable emotions solely by adjusting the emotional dimensions. The alignment between our results and established emotional speech patterns suggests that the model effectively captures the expected emotional characterizations.
Figure 4: XAB test to assess the emotion intelligibility for 4 emotion pairs (‘Surprise’ and ‘Alert’, ‘Relaxed’ and ‘Protected’, ‘Excited’ and ‘Pleasure’, ‘Angry’ and ‘Anxious’). ED values were manually assigned based on Table 1 to simulate different emotions (‘ED Control’).
3.4 Subjective Evaluation
We conducted listening experiments, with 10 native English speakers each listening to a total of 260 synthesized speech samples. We first evaluate our proposed framework on the zero-shot TTS task, where the ED is predicted based on the prompt speech (‘Proposed’). We selected ‘CosyVoice’ [13] as the baseline. A mean opinion score (MOS) test is conducted to evaluate the speech naturalness, where participants were asked to score each sample on a 5-point scale (5: Excellent, 4: Good, 3: Satisfactory, 2: Unsatisfactory, 1: Bad). As shown in Table 2, our proposed framework with ED prediction achieves a score of 4.54, outperforming the baseline score of 4.36, while lower than the ground-truth score of 4.80. It shows that our proposed framework with ED prediction achieves better naturalness than the baseline CosyVoice.
We additionally conducted a mean opinion score to assess emotional intelligibility (E-MOS), where participants were asked to rate each speech sample in terms of emotion perception. We evaluated our proposed framework against two baselines: MixedEmotion [25] and CosyVoice [13]. As shown in Fig.3, our proposed framework with ED prediction consistently outperforms CosyVoice with a higher E-MOS score, while MixedEmotion, which trains on in-domain emotional data, achieves the top score below the ground truth. In contrast, both CosyVoice and our proposed system rely solely on zero-shot emotion cloning. These results indicate that our proposed method effectively captures and replicates emotional nuances in synthesized speech, thereby enhancing emotion cloning performance.
We further conducted four XAB tests, grouping 4 emotion pairs 2 : 1) ‘Anger’ vs. ‘Anxious’; 2) ‘Excited’ vs. ‘Pleasure’; 3) ‘Protected’ vs. ‘Relaxed’ ; 4) ‘Alert’ vs. ‘Surprise’. For each pair, participants were asked to correctly match the emotions, such as choosing between option A: X is Anger and Y is Anxiety and option B: X is Anxiety and Y is Anger. The results are summarized in Fig. 4. For all 4 emotion pairs, the correct match consistently exceeds 50%, with the ‘Anger vs. Anxious’ pair achieving the highest accuracy at around 83.75%. The results in Fig.4 demonstrate a positive alignment with 2 key findings in [18]:
  • ‘Dominance’ is a key factor in distinguishing ‘Angry’ from ‘Anxious’, ‘Alert’ from ‘Surprised’, and ‘Relaxed’ from ‘Protected’ [^4] The first emotion in each pair involves dominance; the second emotion involves submissiveness.;
  • ‘Pleasure’ and ‘Arousal’ combine to contribute to ‘Excitement’;
These observations indicate that our proposed TTS system can render emotions in accordance with psychological theories on the multidimensional nature of emotions, incorporating factors such as pleasure, arousal, and dominance. This enables our model to synthesize speech with a variety of emotional styles, enhancing both the naturalness and emotional intelligibility, and effectively bridging psychological theory with speech synthesis applications.
4 Conclusion
In this paper, we proposed an emotional TTS framework that integrates established emotion theory with language model-based speech generation. By controlling pleasure, arousal, and dominance, our proposed system synthesizes a wide range of emotional speech without relying on large-scale emotion-labeled data, while preserving naturalness in zero-shot scenarios. This work highlights the value of combining psychological insights with data-driven modeling. Future directions include exploring dynamic within-utterance emotion control and multilingual adaptation, paving the way for more versatile, affect-aware TTS solutions.